the total nnumber of words that can be made by writing the letters of the word PARAMETER SO THAT NO VOWEL IS BETWEEN 2 CONSONANTS

Grade:12

4 Answers

Sudheesh Singanamalla
13 years ago

Dear Shivam ,

the given word is PARAMETER.

and between every 2 consonants there is a vowel.

so given condition is that THERE IS NO VOWEL BETWEEN 2 CONSONANTS

so for that , we have to take all the vowels to ONE SIDE and FOR NOW treat them as a SINGLE letter.

vowels are : A , A , E , E

Consonants : P , R , M , T , R

let us take AAEE as one.

so total number = 5 + 1 = 6 (5 consonants + 1 vowel)

and number of words that can be made = 6! = 720

but the vowels can also be arranged in 4! / 2!*2! ways because 'A' appears twice and 'E' appears twice. =  6

so total number of words = 720 * 6 = 4320.

 

Please approve !

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Keep all vowels together and all consonants together we have following ways:

4! * 4! /2! 2! 2!

 

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

Sandesh
6 years ago
This word parameter  consists 5 consonants and 4 vowels 
Let 4 vowels be x
Hence total no of words =6!/2!
But vowels can arranged among themselves in 4!/(2!×2! )Wqys 
Hence total no of words = 600×6
= 1800 
Atharva
6 years ago
If we take all the vowels together that is AAEE then we can arrange them by 4!/2!*2! = 6 Now all the continents can be arranged by 5!/2!=60And now you can place the consonants in 5 days around the vowels •A•A•E•E• ...... (• denotes the ways you can arrange number of words that can be made by using all the letters of word parameter so that no one is between two consonants)So the total answer is 6*60*5 = 1,800Show the final answer is ,1800